Use the substitution method to solve the linear system.
y = x - 4
4x+y = 26

Answer by checkley75(3666)

Y=X-4
4X+(X-4)=26
4X+X-4=26
5X=26+4
5X=30
X=30/5
X=6 ANSWER.
Y=6-4
Y=2 ANSWER.
PROOF
4*6+2=26
24+2=26
26=26